A RECTANGULAR PARK WAS BUILT SO THAT ITS LENGTH IS 3 TIMES ITS WIDTH. THE PERIMETER OF THE PARK IS 24 YARDS. WHAT ARE TEH WIDTH AND LENGTH OF THE PARK?
Answer by checkley77(12844)

L=3W
2L+2W=P
2*3W+2W=24
6W+2W=24
8W=24
W=24/8
W=3 YDS. FOR THE WIDTH.
L=3*3=9 YDS. FOR THE LENGTH.
PROOF:
2*9+2*3=24
18+6=24
24=24